# No Language Data Left Behind: A Comparative Study of CJK Language Datasets in the Hugging Face Ecosystem

### v1

@@ Abstract
Recent advances in Natural Language Processing (NLP) have underscored the crucial role of high-quality datasets in building large language models (LLMs). However, while extensive resources and analyses exist for English, the landscape for East Asian languages—particularly Chinese, Japanese, and Korean (CJK)—remains fragmented and underexplored, despite these languages together serving over 1.6 billion speakers. To address this gap, we investigate the HuggingFace ecosystem from a cross-linguistic perspective, focusing on how cultural norms, research environments, and institutional practices shape dataset availability and quality. Drawing on more than 3,300 datasets, we employ quantitative and qualitative methods to examine how these factors drive distinct creation and curation patterns across Chinese, Japanese, and Korean NLP communities. Our findings highlight the large-scale and often institution-driven nature of Chinese datasets, grassroots community-led development in Korean NLP, and an entertainment- and subculture-focused emphasis on Japanese collections. By uncovering these patterns, we reveal practical strategies for enhancing dataset documentation, licensing clarity, and cross-lingual resource sharing—ultimately guiding more effective and culturally attuned LLM development in East Asia. We conclude by discussing best practices for future dataset curation and collaboration, aiming to strengthen resource development across all three languages.
@@ 1 Introduction
While CJK languages play an increasingly important role in global NLP research, several critical issues need to be addressed. First, there is limited understanding of how dataset creation patterns differ across these language communities—and how those differences reflect their respective NLP ecosystems. Second, although cultural and institutional factors evidently influence dataset generation, their specific impacts on dataset characteristics and quality have yet to be systematically investigated. Third, the potential for cross-lingual synergies among CJK languages remains largely untapped, despite their many shared cultural and linguistic foundations. For instance, while benchmarks like MMLUHendrycks et al. (2020) are being generated in multiple languages, insufficient comparative analysis exists regarding how these variants differ from traditional parallel corpora or how download patterns across regions reflect distinct cultural preferences. To address these challenges, this study investigates the HuggingFace ecosystem for Chinese, Japanese, and Korean NLP resources, focusing on dataset development and usage patterns. The specific objectives of this work are to:
@@ 2 Related Work
Documentation efforts for CJK language resources have expanded progressively over time. For Chinese, Tao et al. (2009) introduced systematic approaches for constructing and evaluating linguistic data resources. More recently, Li et al. (2023a) compiled a comprehensive repository of Chinese datasets, emphasizing detailed accessibility and classification frameworks. In the Korean context, Cho et al. (2020) provided a thorough survey of existing NLP resources, stressing the need for enhanced documentation and broader data accessibility. Building on this, Cho et al. (2023) examined the historical interplay between dataset creation and NLP model development, underscoring how local research cultures influence resource development. However, these earlier studies largely predate the rise of large generative models (e.g., GPT) and relied on GitHub as the primary hub for open-source NLP efforts. Since then, new platforms such as HuggingFace have substantially reshaped data-sharing practices, yet systematic, cross-lingual analyses of CJK datasets within these newer ecosystems remain rare.
In parallel, major multilingual projects— including BigScience(Le Scao et al., 2023), CC100 (Wenzek et al., 2019), and LAION (Schuhmann et al., 2022)—have expanded the global availability of non-English data. While these initiatives contribute valuable resources, they do not investigate the deeper cultural or linguistic factors affecting dataset usage within specific communities such as CJK. Moreover, the broader movement toward improved dataset documentation (e.g., Datasheets for Datasets; Gebru et al., 2021) suggests increasing awareness of ethical and practical considerations, yet how such guidelines are adopted varies widely across languages and platforms.
Over the past few years, HuggingFace has emerged as a central repository for sharing models, datasets, and evaluation metrics in NLP. Studies focusing on HuggingFace have flagged challenges such as inconsistent documentation (Yang et al., 2024), limited transparency about training data (Pepe et al., 2024), and ambiguous licensing information. These issues underscore the need to examine dataset curation and documentation in specific cultural contexts, as local norms and priorities can shape how resources are developed, annotated, and shared (Lhoest et al., 2021). Additionally, recent work by Darg̀is et al. (2024) offers valuable insights into building systematic evaluation frameworks for languages with particular linguistic and cultural traits—an approach that could inform the future assessment of CJK datasets. Nonetheless, no existing study has comprehensively analyzed how the unique ecosystem of Chinese, Japanese, and Korean language datasets on HuggingFace is shaped by local research cultures, licensing preferences, and community-driven development.
Building on these perspectives, our research provides the first large-scale, comparative analysis of CJK datasets within the HuggingFace ecosystem. We investigate how cultural contexts, platform-specific practices, and documentation standards influence dataset usage patterns and resource quality. Additionally, we highlight potential synergies for cross-lingual collaboration and offer insights that may guide more effective and ethically grounded dataset curation for East Asian NLP.
Table 1: Analysis framework for CJK datasets, organized by category.
@@ 3.1 Data Collection
We systematically collected dataset information from the HuggingFace platform using their Datasets API [^1] Datasets were retrieved from https://huggingface.co/datasets using language filters and download frequency sorting. All statistics were recorded on January 28, 2025.. Our data collection strategy focused on identifying actively used datasets for each target language (Chinese, Japanese, Korean, and English as a reference).
@@
For each dataset, we extracted metadata fields, which we organized into four main categories reflecting our analysis focus:
  • Scale & Composition: Dataset size, language combinations, task types, and temporal release patterns
  • Development Patterns: Ownership information, licensing details, and community activity metrics
  • Documentation Quality: Dataset cards, README files, academic citations, and metadata completeness
  • Cultural Characteristics: Domain focus, resource development approaches, and community-specific patterns
In addition, we collected the complete dataset cards to analyze documentation practices and cultural characteristics in depth. For research transparency and reproducibility, we release the full metadata and dataset card contents as a public resource on the Hugging Face Hub: [^2] https://huggingface.co/datasets/Dasool/huggingface-cjk-metadata Dasool/huggingface-cjk-metadata.
@@ 4.1 Overview of CJK Datasets
Table 2: Dataset size distribution across languages.
@@ 4.2 Comparative Characteristics of CJK Datasets
Figure 3 presents the distribution of task categories across languages through a heatmap visualization [^3] Task categories group related NLP tasks based on their functionality (e.g., Text Generation includes text-generation, language-modeling, fill-mask; Question Answering includes question-answering, multiple-choice, extractive-qa).. Text Generation emerges as the dominant task across all languages, with particularly high proportions in Chinese (39.3%), Japanese (37.8%), and Korean (36.0%) datasets compared to English (23.2%). Question Answering follows as the second most common task, with moderate variations: Korean (30.9%) and English (29.7%) show higher proportions than Chinese (21.7%) and Japanese (26.9%).
@@
Analysis of instruction-tuning datasets [^4] We identify instruction datasets through ‘instruct’ keywords and common dataset names in metadata. reveals distinct patterns across languages. Among the top 700 most downloaded datasets for each language, instruction-tuning datasets show notable presence: Korean (13.3%), Japanese (12.6%), Chinese (12.3%), and English (7.0%). This higher proportion in CJK languages compared to English suggests particularly active instruction-tuning development in these communities.
Analysis of temporal patterns (Figure 4) reveals distinct characteristics across languages. Chinese datasets show notable early activity in 2023, with peaks around 7 releases. Korean datasets demonstrate the most dramatic fluctuations in 2024, reaching their highest peaks of 12 and 14 releases in mid-2024. Japanese datasets show relatively moderate activity initially but become more active during 2024, reaching peaks of around 7-9 releases. English datasets maintain the most stable pattern throughout the period, typically with 1-4 releases per month.
@@
Figure 5 shows how each language community approaches data licensing [^5] Licenses were classified into five categories: Permissive (e.g., Apache, MIT, CC-BY), Public Domain (e.g., CC0, PDDL), Copyleft/ShareAlike (e.g., GPL, CC-BY-SA), NonCommercial/ND (CC-BY-NC, CC-BY-ND), and Other. For visualization, Permissive and Public Domain categories were combined, as were Copyleft/ShareAlike and NonCommercial/ND.. Two major observations emerge. First, Chinese (53.1%) and English (51.8%) exhibit the highest proportions of Permissive or Public Domain licenses, indicating a shared culture of open access. Both also include moderate segments of Unknown (Chinese 20.3%, English 23.5%) and Copyleft/NonCommercial (19.6% and 13.0% respectively), suggesting a balance between openness and controlled usage.
@@
Figure 6 illustrates three key metrics related to dataset documentation. For academic grounding, measured by the presence of associated arXiv papers, English datasets lead with 39.20%, followed by Chinese (16.97%), while Japanese (10.48%) and Korean (8.89%) trail behind. Regarding structural completeness, measured by the presence of all five recommended Hugging Face dataset card sections, English again leads (18.41%), with Chinese (7.67%), Japanese (1.64%), and Korean (2.46%) showing lower completeness  [^6] HuggingFace dataset card sections: Description, Structure, Creation, Usage, and Additional Info. In terms of documentation depth, measured by average word count, English maintains the highest average (678 words), Chinese stands at 354, and Japanese and Korean remain lower at 168 and 182 respectively. Taken together, these figures indicate a consistent trend: English datasets demonstrate the most thorough and standardized documentation practices, Chinese resources show moderate completeness and Japanese and Korean documentation remains comparatively succinct or under-documented.
@@
Figure 7 illustrates the proportions of datasets contributed by individual and community contributors, research institutes, and companies in the Korean, Chinese, and Japanese communities [^7] Analysis limited to CJK datasets with verifiable institutional affiliations.. The classification was performed through manual verification by the authors, who examined each dataset’s contributor profiles. In all three languages, individual and community contributors dominate overall: Korean datasets lead with 79.2% contributed by individuals and smaller communities, followed by Japanese at 71.9% and Chinese at 61.6%. Research institutes play a stronger role in China (27.3%) than in Korea (11.1%) or Japan (21.0%), suggesting a more prominent institutional involvement in Chinese NLP resource development. Company contributions remain the smallest category across all three, though slightly higher in Chinese (11.1%) compared to Korean (9.7%) and Japanese (7.2%). This distribution highlights the key role of grassroots efforts as primary drivers of dataset creation, while research institutes and companies offer valuable yet comparatively smaller contributions.
@@ 4.3.1 Chinese Dataset Ecosystem
Chinese datasets also demonstrate significant depth in specialized fields. The FreedomIntelligence/CMB Xidong Wang* (2023) collection features a hierarchical structure (6 main categories, 28 subcategories) spanning 11,200 medical questions, thus enabling targeted evaluations in clinical question-answering. Similarly, TCMLM/TCM_Humanities Kang (2024) focuses on Traditional Chinese Medicine, integrating resources from professional certification materials and historical texts.
@@ 4.3.2 Korean Dataset Ecosystem
A unique strength of the Korean dataset ecosystem is its emphasis on content moderation, encompassing hate-speech detection, toxicity filtering, and profanity masking. Popular resources such as jeanlee/kmhas-korean-hate-speech Lee et al. (2022), humane-lab/K-HATERS Lab (2025) and Dasool/KoMultiText Choi et al. (2023) reflect heightened community and research interest in combating harmful or discriminatory language online. However, these moderation-oriented resources raise broader ethical and regulatory questions, such as defining thresholds for hate speech or handling user privacy. Although the Hugging Face platform provides general community guidelines, more detailed policies—particularly for age-restricted or sensitive data—would help standardize safe usage of these potentially sensitive resources. [^8] As of January 2025, based on Hugging Face’s publicly available platform policies and community guidelines.
@@ 4.3.3 Japanese Dataset Ecosystem
Japanese NLP datasets often integrate subcultural or entertainment-related material, an approach that distinguishes them from other CJK resources. For instance, joujiboi/japanese-anime-speech joujiboi (2024) targets automatic speech recognition in anime content, attracting high download counts and demonstrating direct utility for real-world applications such as subtitle generation. Additionally, YANS-official/ogiri-test-with-references YANS-official (2023) captures the distinctive Ogiri comedy tradition, illustrating Japan’s unique comedy culture through multimodal data (text and images). While such resources enrich models’ ability to handle colloquial or creative contexts, they also require careful documentation of stylistic nuances and potential copyright constraints. Many subcultural datasets involve fan works or licensed content, which often precludes fully open licenses. Researchers must therefore verify these constraints to avoid unintended usage restrictions or downstream complications.
@@ 5 Discussion
Our analysis of Chinese, Japanese, and Korean (CJK) datasets on Hugging Face reveals key trends for LLM development in East Asia. Each language community exhibits distinct characteristics and challenges: Chinese datasets show strong institutional backing but inconsistent documentation; Korean datasets demonstrate active community participation but face licensing and documentation issues; and Japanese datasets feature unique subcultural content while dealing with copyright constraints and synthetic data quality concerns.
From a practical standpoint, three points warrant attention. First, the diversity of licensing schemes—particularly the prevalence of “Unknown” licenses in Korean datasets and copyleft or restricted licenses in Japanese resources—complicates commercial or collaborative projects. Although Hugging Face offers an official dataset card template, many CJK datasets only partially adopt its recommended sections. More consistent adherence to existing templates or integration of additional guidelines (e.g., “Datasheets for Datasets” Gebru et al. (2021)) could enhance clarity and reusability. Second, domain clustering (e.g., medical for Chinese, subcultural for Japanese, moderation tasks for Korean) may leave other areas underserved. Broadening coverage could yield more general-purpose LLM development. Third, culturally specific content—dialects, humor traditions, and local censorship norms—requires transparent documentation, given that simple translations often miss nuanced cultural meanings.
Despite these differences, there is strong potential for synergy. Joint benchmarks spanning all three languages could facilitate fairer cross-lingual performance comparisons, while unified documentation frameworks could standardize metadata, licensing, and annotation. Collaborative consortia could streamline data-cleaning pipelines and repository best practices. Finally, future work may integrate linguistic experts for dataset reviews, conduct controlled experiments to gauge the effects of domain specialization, and broaden the platform scope beyond Hugging Face. Overall, our findings underscore both the richness and fragmentation of CJK resources, suggesting that clearer licensing practices, documentation standards, and cross-lingual collaboration can foster a robust and inclusive ecosystem for LLM across East Asia.

### v2

# No Language Data Left Behind: A Cross-Cultural Study of CJK Language Datasets in the Hugging Face Ecosystem

@@ Abstract
Recent advances in Natural Language Processing (NLP) have underscored the crucial role of high-quality datasets in building large language models (LLMs). However, while extensive resources and analyses exist for English, the landscape for East Asian languages, particularly Chinese, Japanese, and Korean (CJK), remains fragmented and underexplored, despite these languages serving over 1.6 billion speakers. To address this gap, we investigate the HuggingFace ecosystem from a cross-linguistic perspective, focusing on how cultural norms, research environments, and institutional practices shape dataset availability and quality. Drawing on more than 3,300 datasets, we employ quantitative and qualitative methods to examine how these factors drive distinct creation and curation patterns across Chinese, Japanese, and Korean NLP communities. Our findings highlight the large-scale and often institution-driven nature of Chinese datasets, grassroots community-led development in Korean NLP, and an entertainment and subculture-focused emphasis on Japanese collections. By uncovering these patterns, we reveal practical strategies for enhancing dataset documentation, licensing clarity, and cross-lingual resource sharing, guiding more effective and culturally attuned LLM development in East Asia. We conclude by discussing best practices for future dataset curation and collaboration, aiming to strengthen resource development across all three languages.
@@ 1 Introduction
While CJK languages play an increasingly important role in global NLP research, several critical issues need to be addressed. First, there is a limited understanding of how dataset creation patterns differ across these language communities—and how those differences reflect their respective NLP ecosystems. Second, although cultural and institutional factors evidently influence dataset generation, their specific impacts on dataset characteristics and quality have yet to be systematically investigated. Third, the potential for cross-lingual synergies among CJK languages remains largely untapped, despite their many shared cultural and linguistic foundations. For instance, while benchmarks like MMLUHendrycks et al. (2020) are being generated in multiple languages, insufficient comparative analysis exists regarding how these variants differ from traditional parallel corpora or how download patterns across regions reflect distinct cultural preferences. To address these challenges, this study investigates the HuggingFace ecosystem for Chinese, Japanese, and Korean NLP resources, focusing on dataset development and usage patterns. The specific objectives of this work are to:
@@ 2 Related Work
Documentation efforts for CJK language resources have expanded over time. For Chinese, Tao et al. (2009) introduce systematic approaches for constructing and evaluating linguistic data resources, while Li et al. (2023a) compile a comprehensive repository emphasizing accessibility and classification frameworks. In the Korean context, Cho et al. (2020) survey existing NLP resources, and Cho et al. (2023) examine how local research cultures influence resource development. However, these studies largely predate large generative models and rely on GitHub as the primary hub, while systematic analyses of CJK datasets within newer ecosystems like HuggingFace remain rare.
Major multilingual projects including BigScience(Le Scao et al., 2023), CC100 (Wenzek et al., 2019), and LAION (Schuhmann et al., 2022) expand non-English data availability, yet don’t investigate cultural factors affecting dataset usage within communities. The movement toward improved dataset documentation (Gebru et al., 2021) highlights ethical considerations, though adoption varies across languages and platforms.
HuggingFace has emerged as a central repository for NLP resources, but studies flag challenges including inconsistent documentation (Yang et al., 2024), limited transparency (Pepe et al., 2024), and ambiguous licensing. These issues underscore the need to examine dataset curation in cultural contexts (Lhoest et al., 2021). Recent work by Darg̀is et al. (2024) offers insights for building evaluation frameworks for languages with particular traits, yet no study comprehensively analyzes how CJK dataset ecosystems are shaped by local research cultures, licensing preferences, and community-driven development.
Building on these perspectives, our research provides the first large-scale, comparative analysis of CJK datasets within the HuggingFace ecosystem, investigating how cultural contexts and documentation standards influence dataset usage patterns and resource quality.
@@
Table 1: Analysis framework for CJK datasets, organized by category.
@@ 3.1 Data Collection
We systematically collected dataset information from the HuggingFace platform using their Datasets APIDatasets were retrieved from https://huggingface.co/datasets using language filters and download frequency sorting. All statistics were recorded on January 28, 2025.. Our data collection strategy focused on identifying actively used datasets for each target language (Chinese, Japanese, Korean, and English as a reference).
@@
For each dataset, we extracted metadata across four categories: Scale & Composition (dataset size, language combinations, task types, temporal patterns), Development Patterns (ownership, licensing, community metrics), Documentation Quality (dataset cards, citations, README files, metadata completeness), and Cultural Characteristics (domain focus, development approaches, community patterns).
In addition, we collected the complete dataset cards to analyze documentation practices and cultural characteristics in depth. We will release the full metadata and dataset card contents as a public resource upon publication.
@@ 4.1 Overview of CJK Datasets
Table 2: Dataset size distribution across languages.
@@ 4.2 Comparative Characteristics of CJK Datasets
Figure 3 presents the distribution of task categories across languages through a heatmap visualizationTask categories group related NLP tasks based on their functionality (e.g., Text Generation includes text-generation, language-modeling, fill-mask; Question Answering includes question-answering, multiple-choice, extractive-qa).. Text Generation emerges as the dominant task across all languages, with particularly high proportions in Chinese (39.3%), Japanese (37.8%), and Korean (36.0%) datasets compared to English (23.2%). Question Answering follows as the second most common task, with moderate variations: Korean (30.9%) and English (29.7%) show higher proportions than Chinese (21.7%) and Japanese (26.9%).
@@
Analysis of instruction-tuning datasetsWe identify instruction datasets through ‘instruct’ keywords and common dataset names in metadata. reveals distinct patterns across languages. Among the top 700 most downloaded datasets for each language, instruction-tuning datasets show notable presence: Korean (13.3%), Japanese (12.6%), Chinese (12.3%), and English (7.0%). This higher proportion in CJK languages compared to English suggests particularly active instruction-tuning development in these communities.
Analysis of temporal patterns (Figure 5) reveals distinct characteristics across languages. Chinese datasets show notable early activity in 2023, peaking around 7 releases. Korean datasets demonstrate dramatic fluctuations in 2024, reaching highest peaks of 12-14 releases mid-2024. Japanese datasets show moderate initial activity but increased activity during 2024, reaching peaks of 7-9 releases. English datasets maintain stable patterns throughout, typically with 1-4 releases monthly.
@@
Figure 4 shows how each language community approaches data licensingLicenses were classified into five categories: Permissive (e.g., Apache, MIT, CC-BY), Public Domain (e.g., CC0, PDDL), Copyleft/ShareAlike (e.g., GPL, CC-BY-SA), NonCommercial/ND (CC-BY-NC, CC-BY-ND), and Other. For visualization, Permissive and Public Domain categories were combined, as were Copyleft/ShareAlike and NonCommercial/ND.. Two major observations emerge. First, Chinese (53.1%) and English (51.8%) exhibit the highest proportions of Permissive or Public Domain licenses, indicating a shared culture of open access. Both also include moderate segments of Unknown (Chinese 20.3%, English 23.5%) and Copyleft/NonCommercial (19.6% and 13.0% respectively), suggesting a balance between openness and controlled usage.
@@
Figure 6 illustrates three key metrics related to dataset documentation. For academic grounding, measured by the presence of associated arXiv papers, English datasets lead with 39.20%, followed by Chinese (16.97%), while Japanese (10.48%) and Korean (8.89%) trail behind. Regarding structural completeness, measured by the presence of all five recommended Hugging Face dataset card sections, English again leads (18.41%), with Chinese (7.67%), Japanese (1.64%), and Korean (2.46%) showing lower completeness HuggingFace dataset card sections: Description, Structure, Creation, Usage, and Additional Info. In terms of documentation depth, measured by average word count, English maintains the highest average (678 words), Chinese stands at 354, and Japanese and Korean remain lower at 168 and 182 respectively. Taken together, these figures indicate a consistent trend: English datasets demonstrate the most thorough and standardized documentation practices, Chinese resources show moderate completeness and Japanese and Korean documentation remains comparatively succinct or under-documented.
@@
Figure 7 illustrates the proportions of datasets contributed by individual and community contributors, research institutes, and companies in the Korean, Chinese, and Japanese communitiesAnalysis limited to CJK datasets with verifiable institutional affiliations.. Individual and community contributors dominate across all three languages: Korean datasets lead with 79.2%, followed by Japanese at 71.9% and Chinese at 61.6%. Research institutes play a stronger role in China (27.3%) than in Korea (11.1%) or Japan (21.0%), suggesting more prominent institutional involvement in Chinese NLP resource development. Company contributions remain the smallest category across all three, though slightly higher in Chinese (11.1%) compared to Korean (9.7%) and Japanese (7.2%). This highlights the key role of grassroots efforts as primary drivers of dataset creation.
@@ 4.3.1 Chinese Dataset Ecosystem
Chinese datasets also demonstrate significant depth in specialized fields. The FreedomIntelligence/CMB Wang et al. (2023) collection features a hierarchical structure (6 main categories, 28 subcategories) spanning 11,200 medical questions, thus enabling targeted evaluations in clinical question-answering. Similarly, TCMLM/TCM_Humanities Kang (2024) focuses on Traditional Chinese Medicine, integrating resources from professional certification materials and historical texts.
@@ 4.3.2 Korean Dataset Ecosystem
A unique strength of the Korean dataset ecosystem is its emphasis on content moderation, encompassing hate-speech detection, toxicity filtering, and profanity masking. Popular resources such as jeanlee/kmhas-korean-hate-speech Lee et al. (2022), humane-lab/K-HATERS Lab (2025) and Dasool/KoMultiText Choi et al. (2023) reflect heightened community and research interest in combating harmful or discriminatory language online. However, these moderation-oriented resources raise broader ethical and regulatory questions, such as defining thresholds for hate speech or handling user privacy. Although the Hugging Face platform provides general community guidelines, more detailed policies—particularly for age-restricted or sensitive data—would help standardize safe usage of these potentially sensitive resources.As of January 2025, based on Hugging Face’s publicly available platform policies and community guidelines.
@@ 4.3.3 Japanese Dataset Ecosystem
Japanese NLP datasets often integrate subcultural or entertainment-related material, an approach that distinguishes them from other CJK resources. For instance, joujiboi/japanese-anime-speech joujiboi (2024) targets automatic speech recognition in anime content, attracting high download counts and demonstrating direct utility for real-world applications such as subtitle generation. Additionally, YANS-official/ogiri-test-with-references YANS-official (2023) captures the distinctive Ogiri comedy tradition, illustrating Japan’s unique comedy culture through multimodal data (text and images). While such resources enrich models’ ability to handle colloquial or creative contexts, they also require careful documentation of stylistic nuances and potential copyright constraints. Many subcultural datasets involve fan works or licensed content, which often preclude fully open licenses. Researchers must therefore verify these constraints to avoid unintended usage restrictions or downstream complications.
@@ 5 Discussion
Our analysis reveals distinct characteristics across CJK dataset ecosystems: Chinese datasets show strong institutional backing but inconsistent documentation; Korean datasets demonstrate community-driven development but face licensing gaps; and Japanese datasets emphasize subcultural content while dealing with copyright constraints.
Three practical issues warrant attention. First, licensing diversity (particularly "Unknown" licenses in Korean datasets and restricted licenses in Japanese resources) complicates collaborative projects. More consistent adherence to guidelines like "Datasheets for Datasets" Gebru et al. (2021) could enhance reusability. Second, domain clustering (medical for Chinese, subcultural for Japanese, and moderation for Korean) may underserve other areas needed for general-purpose LLM development. Third, culturally specific content requires transparent documentation, as simple translations miss nuanced cultural meanings.
Despite these differences, strong synergy potential exists. Joint benchmarks could facilitate cross-lingual comparisons, while unified documentation frameworks could standardize metadata and licensing. Our findings underscore both the richness and fragmentation of CJK resources, suggesting that clearer practices and cross-lingual collaboration can foster a robust ecosystem for East Asian LLM development.